A25. What is the molar solubility of cobalt carbonate, CoCO3, in an aqueous solution containing 0.3 M sodium carbonate at 25 °C? (Cobalt carbonate Ksp = 1.0 × 10–10)

(A) 1.0 × 10–5 M

(B) 5.48 × 10–6 M

(C) 0.3 M

(D) 3.3 × 10–10 M

(E) 1.0 × 10–10 M
Answer #1

Na2CO3 here is Strong electrolyte

It will dissociate completely to give [CO32-] = 0.3 M

At equilibrium:

CoCO3 <----> Co2+ + CO32-

   s 0.3 + s

Ksp = [Co2+][CO32-]

1*10^-10=(s)*(0.3+ s)

Since Ksp is small, s can be ignored as compared to 0.3

Above expression thus becomes:

1*10^-10=(s)*(0.3)

1*10^-10= (s) * 0.3

s = 3.333*10^-10 M

Answer: 3.3*10^-10 M